From acd@roe.ac.uk Wed Jul 24 09:16:21 2013 Date: Wed, 24 Jul 2013 08:06:25 +0000 From: Clive Davenhall To: Nigel C. Hambly , Eckhard Sutorius , Nicholas Cross , Rob Blake , Ross Collins , "Read, Mike" , Mark S. Holliman Cc: Lorenzo Rimoldini , Bob Mann , Jim Emerson , ktn keithnoddle.org , Norman Gray , Stephen Warren , Tom Shanks , Simon Dye , Stelios Voutsinas , Tom Kerr , Mike Irwin , Andy Lawrence , Nigel Metcalfe Subject: WFAU VDFS Science Archives progress meeting minutes, 12 July 2012 WFAU VDFS Science Archives progress meeting minutes, 12 July 2013 Present: RSC, RGM, MSH, ETWS, RPB, ACD Apologies: NJC, NCH, MAR, STV, KTN, AL, JPE DoNM: 10am Friday July 26th 2013 in the Villas meeting room MAR in the chair (room is booked). Actions discharged this time: ----------------------------- None. Actions partly discharged but continuing: ----------------------------------------- ACTION: MAR to have a go at releasing other VISTA PI DBs (e.g. for JPE) Continues: MAR was not present, but he was believed to have discussed some of the issues with JPE. Actions carried forward from 14/06/2013 meeting ----------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACD noted that this action was already continuing during his previous sojourn with WFAU over the winter of 2011-12. Specific points and new actions: -------------------------------- Project management: WFAU quarterly was being arranged for late August. The data is not yet fixed, but the most promising alternative is currently 27/8/13, pm. WFCAM, VISTA and VST updates: * Nothing to report. Comments and issues arising from CASU minutes: * Minutes were circulated by NCH on 1/7/13 for the CASU meetings of 29/5/13 and 26/6/13. * ETWS noted that temporary optical baffles had been installed on the VST; it would be interesting to see if they improve the stray light correction. Networking: * Nothing to report. WSA/VSA/OSA/GES Operations: * ETWS reported that the transfer of the v1.3 re-processed data had finished. The metadata is ingested and JPGs have been created as well. Transfer of the latest released data from 04/13 is going on. Will start ingesting of the catalogues next week. * ETWS also reported that there were some new OSA data to transfer. * RPB reported that the VMC P90 release database was currently in preparation. It would include all the 1.2 data. The deep products were currently being created. * ACD noted that the FITS template for the next data release had been received and the data themselves were expected to start arriving next week. * Overall operations were reasonably stable and progressing smoothly. Hardware and Systems: * MSH reported that a new database node had been ordered. He did not know when it would arrive. Initially it is to be formatted using RAID 10 and used for the VVV database. If the RAID 10 configuration did not prove satisfactory it would then be reformatted using the more traditional RAID 5. In the RAID 10 configuration the node would offer approx. 30 TByte of storage. * Otherwise the hardware situation was satisfactory; everything was running. Documentation had been written and RPB (and others) knew how to access it. * MSH noted that he would on leave shortly. Arrangements to provide cover were in place. Survey Data Release: * No releases within the last two weeks. * ETWS reported that there had been some discussion of ESO UCDs in DR3; he was in communication with Norman Gray on this issue. * There was some discussion about keeping the timing of public releases by WFAU and ESO consistent. ACTION: RPB to investigate maintaining consistency between WFAU and ESO releases. Software: * ETWS reported that he had: - Rewritten CU1 to change the singular day multi-threaded transfer to a multiple day multi-threaded transfer. This enhances transfer speeds from 30-35MB/s to up to 50MB/s. - Fixed code to better catch files that got truncated due to network hick-ups. Also included automated fixes for non-standard FITS files containing lower case 'e' exponential numbers. - Fixed CU2 to not leave zombie processes lying around while running, which might have contributed to the kernel panic on neferefre. ['cos there is nothing like a zombie for causing a panic - ACD]. * RSC reported that he and RPB were making performance improvements to CU6. The driver for this work to was to improve the performance until it was fast enough to allow processing of VVV DR2 to be completed in the time available. Non-survey Data Release: * Nothing to report. Astrogrid deployment, VO & Data Analysis services: * MSH noted that there was no specific VO work to report. However he had been working on the MonetDB column-oriented (rather than the more usual row-oriented) database. This work was in conjunction with a visitor working with Malcolm Atkinson. It was progressing. Miscellaneous: * RPB noted that a copy of the GALEX GR7 was on its way: it was currently crossing the Atlantic in a jiffy bag. * RGM noted the arrival of Nathan James Hambly. All offered congratulations. * RGM reported that he would be going to ADASS. He should present a poster about WFAU's work and the existing one needed to be revised. ACTION: All to think about items that could be included in the WFAU poster. ACD, 12/7/13. -- Scanned by iCritical.